4,5,9,18,34,59...........whts after 59
Answer by williamson.family05(2)

To find what comes after 59...You have to find what the sequence is.
4 5 9 18 34 59...
4+1=5, 5+4=9, 9+9=18, 18+16=34, 34+25=59 then 1+4=3, 4+5=9, 9+7=16, 16+9=25
4, 5, 9, 18, 34, 59...

1, 4, 9, 16, 25

3, 5, 7, 9
So the next number in the odd sequence3,5,7,9... would be 11. So 25+11=36.
Then the next number in the sequence 1,4,9,16,25... would be 36. So 59+36=95.
So the number after 59 would be 95.